OpenStudy (anonymous):

Make Y the subject of the formulae below question... the square root of YR=Q... show working please!!!!!!... help me

OpenStudy (ash2326):

Is this your question? \[\sqrt{YR}=Q\]

OpenStudy (anonymous):

yes yyes

OpenStudy (anonymous):

@ash2326

OpenStudy (ash2326):

First Step square both the sides \[(\sqrt{YR})^2=Q^2\] \[YR=Q^2\] Now divide both the sides by R can you do this step @Shanelle ?

OpenStudy (anonymous):

yes i can i will try

OpenStudy (anonymous):

@ash2326

OpenStudy (ash2326):

good, try this and if you are stuck.do Let me know

OpenStudy (anonymous):

kay is the answer y= Q squarre

OpenStudy (ash2326):

You have missed R

OpenStudy (anonymous):

i dnt understand

OpenStudy (anonymous):

@ash2326

OpenStudy (ash2326):

\[YR=Q^2\] divide both sides by R \[Y=\frac {Q^2}{R}\]

OpenStudy (anonymous):

i dont get it @ash2326

OpenStudy (ash2326):

You got till here, ?? \[YR=Q^2\]

OpenStudy (anonymous):

what

OpenStudy (ash2326):

did you understand the first step? \[YR=Q^2\]

OpenStudy (anonymous):

not really go over back datt step plz

OpenStudy (ash2326):

We have \[\sqrt{YR}=Q\] square both sides \[{\sqrt{YR}}^2=Q^2\] we get \[YR=Q^2\]

OpenStudy (ash2326):

do you get this?

OpenStudy (anonymous):

yes bt why u remove the square root sign

OpenStudy (ash2326):

if we square then the root sign gets removed \[(\sqrt 2)^2=2\]

OpenStudy (anonymous):

ok continue

OpenStudy (anonymous):

pl

OpenStudy (anonymous):

please

OpenStudy (ash2326):

Next step is to divide both sides by R YR/R=Q^2/R R on the left side will get cancelled, We get finally Y= Q^2/R

OpenStudy (anonymous):

me lost @ash2326

OpenStudy (ash2326):



OpenStudy (ash2326):



OpenStudy (anonymous):

But i asked you b4 if tht was the answer and u didnt reply

OpenStudy (anonymous):

but i undersstan now

OpenStudy (anonymous):

tthanks for you help

OpenStudy (ash2326):

welcome:)

OpenStudy (anonymous):

Express the following in terms of M........ the square root of MN=W